OpenStudy (anonymous):

Find the indefinite integral of integral of [(-x)/( (x+1) - (sqrt of x+1)]dx

OpenStudy (anonymous):

\[\int\limits_{?}^{?}-x / [(x+1) - (\sqrt{x+1})] dx\]

OpenStudy (anonymous):

does the equation make sense teh way i wrote it

OpenStudy (anonymous):

its an indefinite integral not a definite one so tahts y there is still ? on teh top n side

OpenStudy (anonymous):

\[\int\frac{-x}{x+1}dx\ - \int \sqrt{x+1}dx\] this?

OpenStudy (anonymous):

no u see how u separated the squareroot of x+1....that is supposed to b included in teh denominator...so its teh integral of num: -x and den: (x+1) - (sqrt of x+1)

OpenStudy (anonymous):

i mean how did u do write it liek a fraction..wen i do it its always s slash sign

OpenStudy (anonymous):

\[\int \frac{-x dx}{(x+1) - \sqrt{x+1}}\] this?

OpenStudy (anonymous):

yes exxcept the dx is on the side of teh whole equation instead of in teh numerator...does taht make a dffnrce

OpenStudy (anonymous):

it does not

OpenStudy (anonymous):

ok so what would you do first?

OpenStudy (anonymous):

well we are supposed to see if any of the der look like a derivative of anotehr so we can set taht to u

OpenStudy (anonymous):

u-substitution..wil it help here?

OpenStudy (anonymous):

you tell me, can you do a u sub here? or can you do some algebra to manipulate the problem?

OpenStudy (anonymous):

i guess algebra since i dont see how any of them looks like a derivative of another

OpenStudy (anonymous):

good, so what algebra?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

but see the instructions say to use u-sub though but idk how to do that??

OpenStudy (anonymous):

im not sure how alegbra helps

OpenStudy (anonymous):

yes u sub will come after algebra. ill give you a hint, you can multiply the top and bottom by the same thing and it will be like multiplying the whole thing by 1

OpenStudy (anonymous):

u mean mltiply it by its conjugate

OpenStudy (anonymous):

like multiply the top by (x+1) +(sqrt of x+1)

OpenStudy (anonymous):

is taht wat u mean?

OpenStudy (anonymous):

yes, very good, top and bottom remember.

OpenStudy (anonymous):

ok let me wrk this out...

OpenStudy (anonymous):

so the top will be \[(-x^2 -x) + -x \sqrt{x+1}\]

OpenStudy (anonymous):

for the denominator does it \[(x^2+2x+1) - (x+1) = x^2 +x\]

OpenStudy (anonymous):

ok so combine everything

OpenStudy (anonymous):

integral is \int

OpenStudy (anonymous):

\[(-x^2-2x \sqrt{x+1}) / (x^2+x)\] dx

OpenStudy (anonymous):

the top wont combine like that retry

OpenStudy (anonymous):

just dont combine it like you did and try breaking the fraction up into pieces that you think can be integrated

OpenStudy (anonymous):

hmm im so leave teh numerator as (-x^2- x) +-(x*sqrt of x+1)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how about teh dennominator...can it stay as (x^2 +2x+1) - (x+1) or that cud be simplifeid?

OpenStudy (anonymous):

into (x^2 +x) ?

OpenStudy (anonymous):

wat shud i do next?

OpenStudy (anonymous):

keep the denominator simplified here try breaking it into this \[\frac{-x^{2}-x - x \times \sqrt{x+1}}{x^{2}+x}\] into \[\frac{-x^{2}-x}{x^{2}+x} + \frac{-x \sqrt{x+1}}{x^{2}+x}\]

OpenStudy (anonymous):

now see if you can work with this !

OpenStudy (anonymous):

so the integral of teh sum is the sum of its intergrals right?

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

for teh first part..can u=x^2 adn then du = 2xdx so u can get 1/2d = xdx

OpenStudy (anonymous):

im not too sure wat to make u equal to

OpenStudy (anonymous):

what if you factored out a -1 from the top of the 1st part? What do you get?

OpenStudy (anonymous):

then u get (x^2 +x) / (x^2+x)

OpenStudy (anonymous):

thats 1

OpenStudy (anonymous):

multiply by what you factored out right? Can you do that integral?

OpenStudy (anonymous):

wat do u mean

OpenStudy (anonymous):

-1 * 1

OpenStudy (anonymous):

\[\int (-1 \times 1) dx\] for the first integral now how do you do the second?

OpenStudy (anonymous):

so u factor out a 1- again right to get \[-\int\limits_{}^{}x \sqrt{x+1} / (x^2+x)\]

OpenStudy (anonymous):

now what?

OpenStudy (anonymous):

i dont c how anythg wud cancel or simplify

OpenStudy (anonymous):

look hard lots of stuff does!

OpenStudy (anonymous):

oh so wen u factor (x^2+1) taht becomes x(x+1) which is teh same thing under teh radical

OpenStudy (anonymous):

so u can be (x+1)

OpenStudy (anonymous):

just do one thing at a time and write out your steps draw a picture if you have to and post it!

OpenStudy (anonymous):